OpenStudy (anonymous):

What is the volume of a pyramid with base area B = 8 ft2 and height h = 12 ft? A. 24 ft3 B. 32 ft3 C. 64 ft3 D. 96 ft3

563blackghost (563blackghost):

To find the volume you would use the equation.... \(\Huge{V=\frac{Bh}{3}}\) We input what we know... \(\Huge\color{orange}{V=\frac{8 \times 12}{3}}\)

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

thanks!:D

563blackghost (563blackghost):

Np :D

OpenStudy (anonymous):

@563blackghost b?

563blackghost (563blackghost):

Correct ^^

OpenStudy (anonymous):

kk thnx!